OpenStudy (anonymous):

Suppose that a and b are integers, a ≡ 4 (mod 13), and b ≡ 9 (mod 13). Find the integer c with 0 ≤ c ≤ 12 such that a) c ≡ 9a (mod 13).

ganeshie8 (ganeshie8):

\[c \equiv 9a \pmod {13}\] replace \(a\) by \(4\) and reduce

OpenStudy (anonymous):

But the answer will not be right then.

ganeshie8 (ganeshie8):

what do you get after reducing ?

OpenStudy (anonymous):

Need to solve it after.

ganeshie8 (ganeshie8):

hmm

OpenStudy (anonymous):

I did like this. As c = 9a(mod13). a = 4(mod13), put the value then c = (36(mod 13))mod13 then how it solve further

ganeshie8 (ganeshie8):

reduce 36 in mod 13

ganeshie8 (ganeshie8):

36 = 13*2 + 10 so \(c\equiv 9a \equiv 9(4) \equiv 36\equiv 10 \pmod{13}\)

ganeshie8 (ganeshie8):

and 10 is between 0 and 12 so you're done

OpenStudy (anonymous):

value of c is 10, may i right.

OpenStudy (mathmath333):

c = (36(mod 13))mod13 then how it solve further \(c = ((39-3) (mod 13))mod13 \\ c = ((13\times3 -3) (mod 13))mod13 \\ c = ((-3) (mod 13))mod13 \\ c = (10 )~~(mod13) \\ c = (10 )\\ \)